# MicroLad: 2D-to-3D Microstructure Reconstruction and Generation via Latent Diffusion and Score Distillation

### v1

@@ 3.7 Metrics
$u(\mathbf{r})=1\quad\text{on}\quad r_{x}=0,\qquad u(\mathbf{r})=0\quad\text{on}\quad r_{x}=L.$ (49)
@@
This formulation is sensitive to both phase connectivity and tortuosity [60, 61], making $D_{e}$ an indicator of microstructural functionality.
@@ 4 Results
To demonstrate the 2D-to-3D microstructure reconstruction performance on binary microstructures, a 3D carbonate microstructure was reconstructed using MicroLad, as shown in Figure 3. The original 2D carbonate microstructure exhibits spatially distributed binary phases, where black pixels (0) represent the matrix phase and white pixels (1) represent the pore phase, as illustrated in Figure 3(a). By training the LDM on this dataset and performing 2D-to-3D dimensionality expansion, MicroLad generates multiple spatially connected 3D samples (Figure 3(b)). As seen in the randomly selected cross-sectional slices (Figure 3(c))from the reconstructed 3D volumes, the volumes display coherent and well-connected binary structures that visually resemble the original 2D micrographs. The two-point correlation functions of the generated slices also closely match those of the original 2D dataset (Figure 3(d)), indicating that MicroLad effectively preserves the spatial statistics of the microstructure.
@@
To further validate the reconstruction performance of the proposed MicroLad framework, we compare its accuracy on three-phase SOFC microstructures with two existing methods: MPDD [52, 51] and SliceGAN [36]. As shown in Appendix B, the maximum relative error in terms of two-point correlation for MPDD is 12.44%, and for SliceGAN it is 11.01%. In contrast, MicroLad achieves lower errors when $S_{2}$ guidance is applied, demonstrating strong performance in 2D-to-3D microstructure reconstruction with respect to spatial correlation.
@@ 4.2 Inverse-controlled 2D-to-3D generated microstructure samples
Next, property-guided 2D-to-3D generation is performed using the target $D_{e}$, as shown in Figure 7. As the generation progresses toward the specified target, the phase of interest begins to grow and interconnect, forming an effective pathway for diffusion (or conduction)—not merely increasing its volume fraction (Figure 5) or interfacial area (Figure 6). This trend is particularly evident in the case targeting increased $D_{e}^{\text{Pore}}$, as shown in Figure 7(b), where initially disconnected pores (due to their relatively low volume fraction compared to the YSZ and Ni phases) gradually become interconnected, improving the overall diffusivity. The effect becomes even more pronounced when the target diffusivity is set to twice the maximum value observed in the training dataset, as illustrated in Figure C.7. In this case, the pores form a more continuous and interconnected network, enabling more efficient gas diffusion pathways throughout the volume. These results show that the proposed MicroLad framework can not only generate realistic 3D microstructures from 2D inputs, but also embed physics-based property objectives directly into the generative process. This capability enables a deeper analysis of how microstructures may evolve to meet specific functional requirements, thereby advancing our understanding of the complex relationship between morphological features and effective material properties. Additional generated samples for each case, along with the progression of 2D slices during optimization, are provided in Appendix C.
@@ 5.1 Distributional shifts and objective trends in latent space
To assess whether generation with the proposed MicroLad framework genuinely expands the microstructure analysis space, we can analyze the property distribution as done in Section 4.2. Additionally, we can examine the data distribution in the latent space of the LDM. To investigate the behavior of the inverse-controlled generation and the distance between the 2D-to-3D generated data (i.e., 2D slices within the 3D volume) and the original 2D data, we examine their latent embeddings using uniform manifold approximation and projection (UMAP) [62]. A detailed explanation of the latent embedding procedure and UMAP projection is provided in Appendix D.
@@ 5.3 Slice connectivity and spatial coherence
The UMAP-based distances exhibit the similar trend: both the reconstructed and generated volumes show significantly lower inter-slice distances compared to the unconnected 2D slices from the dataset. This observation is visually supported in Figure 10. In Figure 10(a), the independent 2D slices from the dataset appear scattered in the UMAP projection, with no close spatial grouping between consecutive pairs. In contrast, Figure 10(b) illustrates that consecutive slices from MicroLad-generated 3D volumes remain closely clustered in the latent space, suggesting that spatial coherence is effectively preserved across slices. These findings indicate that the proposed MicroLad framework is capable of generating spatially consistent 2D slices within 3D volumes, thereby maintaining the local continuity necessary for both accurate 2D-to-3D microstructure reconstruction and inverse-controlled generation.
@@ 5.4 Computational cost
Since the proposed MicroLad framework employs latent diffusion in a lower-dimensional space, the L-MPDD procedure for generating initial 3D microstructures is highly efficient. Generating a $64\times 64\times 64$ volume takes less than 10 seconds—comparable to generating a single $64\times 64$ image with a batch size of 64 (wall-clock times were measured on a single NVIDIA RTX 4090 GPU). In contrast, the original MPDD framework required over 30 minutes to produce a single 3D sample with comparable accuracy [51]. While dimensionality expansion via L-MPDD is fast, the most computationally intensive component is SDS applied to enforce target objectives. For each objective, the wall-clock time increases approximately linearly with the number of SDS steps. Appendix E presents a parametric study evaluating the effects of SDS learning rate and objective weights on optimization performance. For volume fraction and surface area targets, 500 SDS steps typically require about one minute to generate a single 3D sample. In contrast, SDS targeting effective diffusivity requires approximately two and a half minutes per sample, due to the overhead of computing gradients through the differentiable FEM solver. Across all objectives, using 3000 SDS steps consistently results in low error between generated samples and their respective targets. Thus, the number of SDS steps can be chosen by balancing computational cost with the desired accuracy of inverse generation. Moreover, the run time of property-guided SDS could be further reduced by replacing the FEM solver with a neural network-based surrogate model, which would allow for more efficient gradient computation. Incorporating such surrogates is recommended as a promising direction for future work.
@@ 5.5 Future work
The proposed MicroLad framework presents considerable potential for extension to a broader class of multiphase materials, enabling the study of microstructural variability and its influence on resulting material properties. One promising direction is to leverage the framework for uncertainty quantification by incorporating user-defined microstructural variability, thereby enhancing its utility in data-driven material design. Furthermore, MicroLad can be extended to a wider range of physics-informed generative modeling tasks, including 2D-to-3D microstructure generation under multiphysics loading conditions and complex material behavior. A particularly promising avenue for future research is the integration of advanced differentiable solvers [66] capable of modeling nonlinear constitutive laws. Such an extension would allow the framework to capture a richer spectrum of microstructural phenomena governed by diverse physical mechanisms.
@@
Appendix A Configuration parameters for the models
@@
where $\lambda_{\text{KL}}$ is a weighting factor set to 0.5 in our implementation. The reconstruction quality of the VAE was evaluated on binary and three-phase microstructure datasets using three metrics: mean absolute error (MAE), peak signal-to-noise ratio (PSNR), and structural similarity index (SSIM). Table A.1 reports the results for multiple latent space sizes, showing that the VAE achieves high-fidelity reconstructions across both binary and multiphase datasets, especially when using a $16\times 16$ latent feature map.
After training the VAE, we train a LDM that operates on the compressed latent space to enable generative modeling. Following the DDPM formulation [42], the LDM defines a forward Markov diffusion process that gradually adds Gaussian noise to a clean latent code. The LDM is trained with 1000 diffusion steps using a linear noise schedule $\beta\in[1\times 10^{-4},2\times 10^{-2}]$. During inference, new samples are generated by reversing the diffusion process starting from pure noise in latent space. These are decoded back to pixel space using the pretrained VAE decoder $\mathcal{D}(\cdot)$. The training hyperparameters for both the VAE and LDM are summarized in Tables A.2 and A.3. The encoder–decoder architecture and full latent U-Net pipeline are detailed in Table A.4,
Table A.1: VAE reconstruction performance across different latent space dimensions and microstructure datasets.
@@
Table A.2: Configuration parameters for the VAE.
@@
Table A.3: Configuration parameters for the LDM.
@@
Table A.4: Encoder–latent-space U-Net–decoder pipeline for MicroLad (latent map $4\times 16\times 16$).
@@
Appendix B Comparison of microstructure reconstruction performance with previous methods
To benchmark the reconstruction accuracy of the proposed MicroLad framework, we compare its performance against two established methods: the MPDD approach and SliceGAN. All methods are evaluated on three-phase microstructures using MAE and relative error metrics for two-point correlation function of each material phase. Table B.1 reports the phase-wise reconstruction errors, showing that MicroLad (with $S_{2}$ guidance) achieves superior fidelity across all phases. For completeness, we also include architectural details of the MPDD and SliceGAN baselines in Tables B.2 and B.3, respectively.
Table B.1: Reconstruction accuracy (three-phase case).
@@ Appendix B Comparison of microstructure reconstruction performance with previous methods
Table B.2: U-Net backbone used for MPDD on $64{\times}64$ images.
@@
Table B.3: SliceGAN generator and discriminator for three-phase volumes ($64^{3}$ voxels).
@@
Appendix C Additional inverse 2D-to-3D generation results
Figure C.1: Visualization of the 2D-to-3D generated samples using MicroLad, along with their target objectives: (a) $V_{f}^{\text{Pore}}$, (b) $V_{f}^{\text{YSZ}}$ and (c) $V_{f}^{\text{Ni}}$
Figure C.2: Visualization of the 2D-to-3D generated samples using MicroLad, along with their target objectives: (a) $SA^{\text{Pore}}$, (b) $SA^{\text{YSZ}}$ and (c) $SA^{\text{Ni}}$
Figure C.3: Visualization of the 2D-to-3D generated samples using MicroLad, along with their target objectives: (a) $D_{e}^{\text{Pore}}$, (b) $D_{e}^{\text{YSZ}}$ and (c) $D_{e}^{\text{Ni}}$
Figure C.4: Progression of three orthogonal cross-sectional slices (through the center of each volume) during the 2D-to-3D microstructure generation process guided by a target objective $V_{f}^{\text{Pore}}$.
Figure C.5: Progression of three orthogonal cross-sectional slices (through the center of each volume) during the 2D-to-3D microstructure generation process guided by a target objective $SA^{\text{Pore}}$.
Figure C.6: Progression of three orthogonal cross-sectional slices (through the center of each volume) during the 2D-to-3D microstructure generation process guided by the target objective $D_{e}^{\text{Pore}}$, set to twice the maximum value in the training dataset.
Figure C.7: Inverse 2D-to-3D generation of three-phase microstructures with controlled relative diffusivity (2 times the maximum in training dataset) for each material phase: (a) generated 3D volumes, (b) variation in 2D cross-sectional images (randomly selected slices) as SDS steps increase, and (c) distribution of variables in the training dataset compared to generated samples with target objectives.
Appendix D Latent-space embedding and UMAP projection

### v3

@@ 3.7 Metrics
$u(\mathbf{r_{1}})=1,\qquad u(\mathbf{r_{0}})=0.$ (49)
@@
This formulation is sensitive to both phase connectivity and tortuosity [60, 61], making $D_{e}$ an indicator of microstructural functionality. Additional details on the FEM implementation are provided in Section C.
@@ 4 Results
To demonstrate the 2D-to-3D microstructure reconstruction performance on binary microstructures, a 3D carbonate microstructure was reconstructed using MicroLad, as shown in Figure 3. The original 2D carbonate microstructure exhibits spatially distributed binary phases, where black pixels (0) represent the matrix phase and white pixels (1) represent the pore phase, as illustrated in Figure 3(a). By training the LDM on this dataset and performing 2D-to-3D dimensionality expansion, MicroLad generates multiple spatially connected 3D samples (Figure 3(b)). As seen in the randomly selected cross-sectional slices (Figure 3(c)) from the reconstructed 3D volumes, the volumes display coherent and well-connected binary structures that visually resemble the original 2D micrographs. The two-point correlation functions of the generated slices (along all three orthogonal directions) also closely match those of the original 2D dataset (Figure 3(d)), indicating that MicroLad effectively preserves the spatial statistics of the microstructure.
@@
To further validate the reconstruction performance of the proposed MicroLad framework, we compare its accuracy on three-phase SOFC microstructures with two existing methods: MPDD [52, 51] and SliceGAN [36]. As shown in Appendix D, the maximum relative error in terms of two-point correlation for MPDD is 12.44%, and for SliceGAN it is 11.01%. In contrast, MicroLad achieves lower errors when $S_{2}$ guidance is applied, demonstrating strong performance in 2D-to-3D microstructure reconstruction with respect to spatial correlation.
@@ 4.2 Inverse-controlled 2D-to-3D generated microstructure samples
Next, property-guided 2D-to-3D generation is performed using the target $D_{e}$, as shown in Figure 7. As the generation progresses toward the specified target, the phase of interest begins to grow and interconnect, forming an effective pathway for diffusion (or conduction)—not merely increasing its volume fraction (Figure 5) or interfacial area (Figure 6). This trend is particularly evident in the case targeting increased $D_{e}^{\text{Pore}}$, as shown in Figure 7(b), where initially disconnected pores (due to their relatively low volume fraction compared to the YSZ and Ni phases) gradually become interconnected, improving the overall diffusivity. The effect becomes even more pronounced when the target diffusivity is set to twice the maximum value observed in the training dataset, as illustrated in Figure E.7. In this case, the pores form a more continuous and interconnected network, enabling more efficient gas diffusion pathways throughout the volume. These results show that the proposed MicroLad framework can not only generate realistic 3D microstructures from 2D inputs, but also embed physics-based property objectives directly into the generative process. This capability enables a deeper analysis of how microstructures may evolve to meet specific functional requirements, thereby advancing our understanding of the complex relationship between morphological features and effective material properties. Additional generated samples for each case, along with the progression of 2D slices during optimization, are provided in Appendix E.
@@ 5.1 Distributional shifts and objective trends in latent space
To assess whether generation with the proposed MicroLad framework genuinely expands the microstructure analysis space, we can analyze the property distribution as done in Section 4.2. Additionally, we can examine the data distribution in the latent space of the LDM. To investigate the behavior of the inverse-controlled generation and the distance between the 2D-to-3D generated data (i.e., 2D slices within the 3D volume) and the original 2D data, we examine their latent embeddings using uniform manifold approximation and projection (UMAP) [62]. A detailed explanation of the latent embedding procedure and UMAP projection is provided in Appendix F.
@@
In addition, while the proposed inverse-controlled microstructure generation formulates microstructural descriptors and properties as objectives to facilitate expansion of the microstructure design and analysis space, the descriptors can also be incorporated as constraints within the same framework for practical inverse design scenarios. For example, in manufacturing-aware design problems, descriptors such as particle size, aspect ratio, or phase connectivity may be restricted by processing limitations or functional requirements. This can be expressed as
$\mathbf{x}^{*}=\underset{\mathbf{x}\in\mathcal{X}}{\arg\min}\;\left[\mathcal{L}_{P}(\mathbf{P}(\mathbf{x}),\mathbf{P}^{*})\right]\quad\text{subject to }\mathbf{M}(\mathbf{x})\in\mathcal{C},$ (51)
where $\mathcal{C}$ denotes the feasible set of microstructural descriptors. In such cases, the framework can accommodate these requirements through penalty or barrier terms, ensuring that the generated microstructures not only explore new regimes but also satisfy practical design constraints. It is suggested that future work place greater emphasis on manufacturability constraints and expand the framework toward manufacturability-aware process–structure–property linkages.
@@ 5.3 Slice connectivity and spatial coherence
The UMAP-based distances exhibit a similar trend: both the reconstructed and generated volumes show significantly lower inter-slice distances compared to the unconnected 2D slices from the dataset. This observation is visually supported in Figure 10. In Figure 10(a), the independent 2D slices from the dataset appear scattered in the UMAP projection, with no close spatial grouping between consecutive pairs. In contrast, Figure 10(b) illustrates that consecutive slices from MicroLad-generated 3D volumes remain closely clustered in the latent space, suggesting that spatial coherence is effectively preserved across slices. These findings indicate that the proposed MicroLad framework is capable of generating spatially consistent 2D slices within 3D volumes, thereby maintaining the local continuity necessary for both accurate 2D-to-3D microstructure reconstruction and inverse-controlled generation.
In the meantime, an important point of discussion is that the present 2D-to-3D microstructure reconstruction/generation framework relies on the assumption that spatial connectivity in 3D space can be represented by connectivity along three orthogonal directions (Eq. (5)). The local consistency assumption (Section 3.4) was likewise built on the optimization of latent slices along these three directions. Although LDM with a microstructure encoder (i.e., VAE) was adopted to enable the computationally efficient 2D-to-3D expansion and to generate 2D slices that remain close to one another in the latent embedding space (Table 3 and Figure 10), consistency in other directions in 3D space (e.g., diagonal) [66] is not explicitly guaranteed. While the proposed approach and its underlying assumptions are practical, given that most experimentally observed 2D micrographs are obtained along one axis or a few orthogonal planes, statistical consistency in non-orthogonal directions or planes in 3D space also needs to be evaluated and addressed in future work. Nevertheless, the proposed approach performed well for both binary and three-phase microstructures, and additional examples, including anisotropic cases presented in Appendix G, further demonstrate that the proposed framework can effectively handle a various types of microstructures.
@@ 5.4 Computational cost
Since the proposed MicroLad framework employs latent diffusion in a lower-dimensional space, the L-MPDD procedure for generating initial 3D microstructures is highly efficient. Generating a $64\times 64\times 64$ volume takes less than 10 seconds—comparable to generating a single $64\times 64$ image with a batch size of 64 (wall-clock times were measured on a single NVIDIA RTX 4090 GPU). In contrast, the original MPDD framework required over 30 minutes to produce a single 3D sample with comparable accuracy [51]. While dimensionality expansion via L-MPDD is fast, the most computationally intensive component is SDS applied to enforce target objectives. For each objective, the wall-clock time increases approximately linearly with the number of SDS steps. Appendix H presents a parametric study evaluating the effects of SDS learning rate and objective weights on optimization performance. For volume fraction and surface area targets, 500 SDS steps typically require about one minute to generate a single 3D sample. In contrast, SDS targeting effective diffusivity requires approximately two and a half minutes per sample, due to the overhead of computing gradients through the differentiable FEM solver. Across all objectives, using 3000 SDS steps consistently results in low error between generated samples and their respective targets. Thus, the number of SDS steps can be chosen by balancing computational cost with the desired accuracy of inverse generation.
It is also worth noting that the main purpose of adopting a differentiable model for guided generation is to eliminate the need for pre-labeling data and training conditional models by directly integrating differentiable simulations into the guided generation framework. This formulation can be adapted depending on the application. If a sufficiently large dataset is available (or can be affordably generated) and reducing inference time is critical—perhaps even more important than minimizing training cost—a conditional formulation (e.g., conditional diffusion models with classifier-free guidance [67]) may be more suitable. In such cases, differentiable model–based guidance may no longer be necessary. The limitation of this approach, however, is that the achievable design or property range tends to be restricted by the data-conditioned generation process, whereas guided generation with differentiable models emphasizes optimization while leveraging the diffusion prior to ensure that generation remains within the unconditional data manifold [68, 53]. Therefore, the appropriate formulation should be selected based on the user’s specific objectives and design requirements.
In addition, to further illustrate the computational cost and scalability of the proposed approach, we provided a 2D-to-3D reconstruction and generation case with a larger microstructure size ($256\times 256\times 256$) in Appendix I. The reconstruction of a 3D volume of this size required only 71 s, with a maximum two-point correlation function relative error of 5.23%, demonstrating both the efficiency and accuracy of the proposed framework for larger microstructure size. Meanwhile, the $D_{e}$-guided generation successfully shifted the property distribution toward the target but required 156.1 minutes, primarily because the FEM computation substantially increased the overall computational cost. This efficiency can be further improved by employing more efficient FEM solvers and differentiable neural network–based surrogate modeling approaches [69, 70], which would enable faster and more efficient gradient computation. Incorporating such surrogates represents a promising direction for future work.
@@ 5.5 Future work
The proposed MicroLad framework presents considerable potential for extension to a broader class of multiphase materials, enabling the study of microstructural variability and its influence on resulting material properties. One promising direction is to leverage the framework for uncertainty quantification by incorporating user-defined microstructural variability, thereby enhancing its utility in data-driven material design. Furthermore, MicroLad can be extended to a wider range of physics-informed generative modeling tasks, including 2D-to-3D microstructure generation under multiphysics loading conditions and complex material behavior. A particularly promising avenue for future research is the integration of advanced differentiable solvers [69] capable of modeling nonlinear constitutive laws. Such an extension would allow the framework to capture a richer spectrum of microstructural phenomena governed by diverse physical mechanisms. The incorporation of a wider variety of microstructures, including those with more phases and even continuous-valued features (e.g., crystallographic orientations), is also suggested, as the proposed approach does not intrinsically restrict the input to discrete phases. This is supported by the fact that latent diffusion models have been successfully applied to colored images as well [48]
@@
Appendix A Refinement of 2D slices on orthogonal planes
To handle unexpected noise or deviations that may occur during the 2D-to-3D dimensionality expansion, a refinement step is applied to the reconstructed 3D volume. This process does not involve training a new denoising model. Instead, it utilizes the originally trained VAE to refine the 2D slices. The key idea is that a VAE trained exclusively on clean data implicitly projects inputs back onto the learned manifold. Because the latent space captures only the distribution of clean samples, details inconsistent with the training distribution are not faithfully reconstructed [71, 72, 73]. Therefore, for the generated 3D volumes, this serves as an additional refinement step applied to the slices along the three orthogonal directions, as illustrated in Figure A.1.
Figure A.1: Schematic of the refinement process for 2D slices of a 3D volume using the trained encoder and decoder.
Appendix B Configuration parameters for the models
@@
where $\lambda_{\text{KL}}$ is a weighting factor set to 0.5 in our implementation. The reconstruction quality of the VAE was evaluated on binary and three-phase microstructure datasets using three metrics: mean absolute error (MAE), peak signal-to-noise ratio (PSNR), and structural similarity index (SSIM). Table B.1 reports the results for multiple latent space sizes, showing that the VAE achieves high-fidelity reconstructions across both binary and multiphase datasets, especially when using a $16\times 16$ latent feature map.
After training the VAE, we train a LDM that operates on the compressed latent space to enable generative modeling. Following the DDPM formulation [42], the LDM defines a forward Markov diffusion process that gradually adds Gaussian noise to a clean latent code. The LDM is trained with 1000 diffusion steps using a linear noise schedule $\beta\in[1\times 10^{-4},2\times 10^{-2}]$. During inference, new samples are generated by reversing the diffusion process starting from pure noise in latent space. These are decoded back to pixel space using the pretrained VAE decoder $\mathcal{D}(\cdot)$. The training hyperparameters for both the VAE and LDM are summarized in Tables B.2 and B.3. The encoder–decoder architecture and full latent U-Net pipeline are detailed in Table B.4,
Table B.1: VAE reconstruction performance across different latent space dimensions and microstructure datasets.
@@
Table B.2: Configuration parameters for the VAE.
@@
Table B.3: Configuration parameters for the LDM.
@@
Table B.4: Encoder–latent-space U-Net–decoder pipeline for MicroLad (latent map $4\times 16\times 16$).
@@
Appendix C Implementation of FEM
In this study, a differentiable FEM solver was implemented in PyTorch to compute the effective diffusivity of SOFC microstructures for each material phase. The solver discretizes the microstructure into linear rectangular elements (e.g., $64\times 64$ or $256\times 256$) and solves the steady-state diffusion equation. Dirichlet boundary conditions are applied to simulate transport through each microstructure phase and obtain the relative diffusivity (Figure C.1). The solver assembles a sparse stiffness matrix based on element-wise contributions and solves the resulting linear system using a PyTorch-based linear solver, enabling automatic gradient computation for property-guided microstructure generation.
Figure C.1: Schematic of the FEM simulation for a given microstructure used to obtain the relative diffusivity.
Appendix D Comparison of microstructure reconstruction performance with previous methods
To benchmark the reconstruction accuracy of the proposed MicroLad framework, we compare its performance against two established methods: the MPDD approach and SliceGAN. All methods are evaluated on three-phase microstructures using MAE and relative error metrics for two-point correlation function of each material phase. Table D.1 reports the phase-wise reconstruction errors, showing that MicroLad (with $S_{2}$ guidance) achieves superior fidelity across all phases. For completeness, we also include architectural details of the MPDD and SliceGAN baselines in Tables D.2 and D.3, respectively.
Table D.1: Reconstruction accuracy (three-phase case).
@@ Appendix D Comparison of microstructure reconstruction performance with previous methods
Table D.2: U-Net backbone used for MPDD on $64{\times}64$ images.
@@
Table D.3: SliceGAN generator and discriminator for three-phase volumes ($64^{3}$ voxels).
@@
Appendix E Additional inverse 2D-to-3D generation results
Figure E.1: Visualization of the 2D-to-3D generated samples using MicroLad, along with their target objectives: (a) $V_{f}^{\text{Pore}}$, (b) $V_{f}^{\text{YSZ}}$ and (c) $V_{f}^{\text{Ni}}$
Figure E.2: Visualization of the 2D-to-3D generated samples using MicroLad, along with their target objectives: (a) $SA^{\text{Pore}}$, (b) $SA^{\text{YSZ}}$ and (c) $SA^{\text{Ni}}$
Figure E.3: Visualization of the 2D-to-3D generated samples using MicroLad, along with their target objectives: (a) $D_{e}^{\text{Pore}}$, (b) $D_{e}^{\text{YSZ}}$ and (c) $D_{e}^{\text{Ni}}$
Figure E.4: Progression of three orthogonal cross-sectional slices (through the center of each volume) during the 2D-to-3D microstructure generation process guided by a target objective $V_{f}^{\text{Pore}}$.
Figure E.5: Progression of three orthogonal cross-sectional slices (through the center of each volume) during the 2D-to-3D microstructure generation process guided by a target objective $SA^{\text{Pore}}$.
Figure E.6: Progression of three orthogonal cross-sectional slices (through the center of each volume) during the 2D-to-3D microstructure generation process guided by the target objective $D_{e}^{\text{Pore}}$, set to twice the maximum value in the training dataset.
Figure E.7: Inverse 2D-to-3D generation of three-phase microstructures with controlled relative diffusivity (twice the maximum in training dataset) for each material phase: (a) generated 3D volumes, (b) variation in 2D cross-sectional images (randomly selected slices) as SDS steps increase, and (c) distribution of variables in the training dataset compared to generated samples with target objectives.
Appendix F Latent-space embedding and UMAP projection